I'm having trouble with the hours... PLEASE HELP!!!

Tyler Guitars makes acoustic and electric guitars. It is struggling to determine the profitability of each guitar and decidinSelling Price & Variable Costs per Guitar Acoustic Variable. Cost Selling Price Acoustic Variable Cost: $95 Electric VariableSelling Price & variable Costs per Guitar Acoustic Variable Cost Selling Price Electric Variable Cost Selling Price ElectricSelling Price & Variable Costs per Guitar Acoustic Variable Cost Selling Price Acoustic Selling Price: $225 Electric Variableimage.pngimage.png150 guitars 3 hrs. 125 guitars 100 guitars 2 hrs. 75 guitars 50 guitars 25 guitars Ohrs. Acoustic Electric O guitars Acoustic1. The company is considering adding a second 8-hour shift for 22 days each month, which would increase fixed costs by $9,0001. The company is considering adding a second 8-hour shift for 22 days each month, which would increase fixed costs by $9,000I'm having trouble with the hours... PLEASE HELP!!!
Answer #1

Hours dedicated to the production of each product = Units produced * Machine hours per unit

Hours dedicated for Acoustic guitars = 131 * 2 = 262 hours

Hours dedicated for Electric guitars = 30 * 3 = 90 hours

Total hours available = 2 shifts * 8 hours * 22 days = 352 hours